OpenStudy (anonymous):

x=(the following equation is all UNDER a radical) 2x+24

OpenStudy (anonymous):

You need to get rid of the radical first eg. x = \[\sqrt{x+3}\] get rid of the root (square root) by squaring both sides of the equation: \[x^{2} = (\sqrt{x+2})^{2}\]

OpenStudy (anonymous):

you are then left with \[x^{2} = x+2\]

OpenStudy (anonymous):

After this step put all the terms on the same side : \[x^{2}-x-2=0\] Then solve for x by factoring or using the quadratic equation: by factoring: (x-2)(x+1)=0 Then using the zero factor property (a)(b)=0: (x-2)=0 and (x+1)=0. Solve for x in each term: x=2 and x=-1